On December 31, 2020, BROWN Bank has a 5-year loan receivable with a face value of ₱6,000,000 dated January 1, 2019 due on December 31, 2023. Interest is payable annually every December 31 at 9%. The borrower made the required interest payment on December 31, 2019 but informed the bank that interest accrued for 2020 will be paid together with the principal at maturity. There is a high probability that remaining interest payments will not be paid. The prevailing market rate is 10%. What is the loan impairment loss for 2020? (Round PV factors to 2 decimals) 


1
Expert's answer
2021-04-19T07:36:23-0400

find the amount of the loan payment every year:

9%

"А=\frac{PVA\times r}{(1-\frac{1}{(1+r)^n})}"

"А=\frac{6 000 000\times 0.09}{(1-\frac{1}{(1+0.09)^5})}=1542 554.74"

10%

"А=\frac{6 000 000\times 0.1}{(1-\frac{1}{(1+0.1)^5})}=1 582 784.88"


1 582 784.88-1 542 554.74=40 230.14